OpenStudy (lgbasallote):

Prove that if m is a perfect square and n is a perfect square, then their product is a perfect square as well

OpenStudy (anonymous):

The key here is\[m^2\cdot n^2=(m\cdot n)^2\]

OpenStudy (lgbasallote):

hmm

OpenStudy (anonymous):

nice...

OpenStudy (anonymous):

It has to do with the fact that exponents distribute over multiplication.

OpenStudy (lgbasallote):

why do you square m and n?

OpenStudy (swissgirl):

Suppose m and n are perfect squares then by definition m=x^2 and n=y^2 The product of m*n=x^2*y^2 Which are perfect squres \(\sqrt{x^2*y^2}\)=x*y Hence The product of two perfect squraes is a perfect square

OpenStudy (anonymous):

oops, ummm

OpenStudy (anonymous):

Well maybe what I should have said was \[m=j^2\]\[n=k^2\]\[j^2\cdot k^2=(j\cdot k)^2\]

OpenStudy (lgbasallote):

yes that;s what i thought too @swissgirl and @wio but for some reason...it's not a valid proof...

OpenStudy (anonymous):

maybe they want you to show an intermediate step like: \[j^2\cdot k^2=j\cdot j \cdot k \cdot k = j\cdot k \cdot j \cdot k =(j\cdot k)(j \cdot k)=(j \cdot k)^2 \]

OpenStudy (swissgirl):

Suppose m and n are perfect squares then by definition m=x^2 and n=y^2 The product of m*n=x^2*y^2=(x*y)^2 Which is perfect squares Hence The product of two perfect squraes is a perfect square

OpenStudy (lgbasallote):

ah i got it. i have to substitude jk into x so it will be the definition of a perfect square \[j^2k^2\implies (jk)^2 \implies x^2\]

OpenStudy (swissgirl):

Yessss

OpenStudy (anonymous):

You really don't have to do the \(x\) think though.

OpenStudy (lgbasallote):

well the definition of a perfect square is n = k^2

OpenStudy (anonymous):

I mean it doesn't make the fact that \(j\cdot k \in \mathbb{Z}\)

OpenStudy (lgbasallote):

still..since this is proving that nm is a perfect square...the definition of a perfect square has to be shown..otherwise it's not a valid proof.